OpenStudy (anonymous):

Solve: 10n – n2 = 0 a) n = 0, n = 10 b) n = -10 c) n = 0 d) n = 0, n = -10 e) n = 10

OpenStudy (tommo_lcfc):

10n - n^2 = n(10 - n) = 0 n = 0 and 10

Parth (parthkohli):

n(10 - n) = 0 10 - n = 0/n 10 - n = 0 (if n is not 0) -n = -10 n = 10 n can be 10 and also 0